### v2

@@ 6.2 Relaxing the ex-ante symmetry
The potential outcomes $\mu=(\mu_{1},\cdots,\mu_{n})\in\mathcal{M}=\otimes_{i=1}^{n}\mathcal{M}_{i}$ where $\mathcal{M}_{i}$ is the set such that $\mu_{i}(1,\cdot)\in\mathcal{M}_{1,i},\mu_{i}(0,\cdot)\in\mathcal{M}_{0,i}$. For all $i\in\{1,\cdots,n\}$, $d\in\{0,1\}$, the function classes $\mathcal{M}_{d,i}$ contains all functions satisfying
@@
    $\mu_{i}(d,\cdot)\in[\psi_{Li},\psi_{Ri}],\forall\mu_{i}(d,\cdot)\in\mathcal{M}_{d,i}$ for some $-\infty<\psi_{Li}<\psi_{Ri}<\infty$;
@@
    $\sup_{\mu_{i}(0,\cdot)\in\mathcal{M}_{0,i}}\Big{|}\mu_{i}(0,\mathbf{d})-\mu_{i}(0,\mathbf{0})\Big{|}$ $=\bar{\phi}_{n}\frac{\alpha_{i}}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}\mathbf{d}_{k},$ (17)
    $\sup_{\mu_{i}(1,\cdot)\in\mathcal{M}_{1,i}}\Big{|}\mu_{i}(1,\mathbf{d})-\mu_{i}(1,\mathbf{1})\Big{|}$ $=\bar{\phi}_{n}\frac{\alpha_{i}}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}\Big{(}1-\mathbf{d}_{k}\Big{)},$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\left\{\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}\right\}(1+o(1)),$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}[(\hat{\tau}_{n}(\mathcal{C}_{n})-\tau_{n,\mu})^{2}]=\left\{\bar{\phi}_{n}^{2}b_{n,\alpha}(\mathcal{C}_{n})^{2}+\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}\right\}(1+o(1)).$
Proof of Theorem 6.2.
See Appendix C.11. ∎
It is easy to show that similarly to what discussed in Section 4.1, the optimization program admits a semi-definite relaxation that depends on the values of $\alpha$ and $\bar{\psi}$.
6.3 Trimming estimators
@@
$\sup_{\mu\in\mathcal{M}}\Big{|}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{|}=\left(\bar{\phi}_{n}b_{n}^{w}(\mathcal{C}_{n})+(\psi_{R}-\psi_{L})\sum_{i-1}^{n}\Big{|}\frac{w_{i}}{\|w\|_{1}}-\frac{1}{n}\Big{|}\right)\cdot(1+o(1))$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot\bar{\psi}(1+o(1)).$
Proof of Theorem 6.3.
See Appendix C.12. ∎
Theorem 6.3 characterizes the worst-case bias and variance of the trimming estimator. Following the discussion for the difference in means estimator, this result suggests a weighted objective function
@@
For the difference in means estimator, the objective function $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)$ defined in Theorem 3.3 can be interpreted as the worst-case MSE, up to $1+o(1)$ terms, as implied by Theorem 3.4. Unfortunately, this exact equivalence no longer holds for general trimming estimators because the worst-case bias and variance are achieved by different configurations of potential outcomes. In fact, we can show that the worst-case MSE is given by the maximum of a quadractic function in the vectors $(\mu_{1}(\mathbf{1}),\ldots,\mu_{n}(\mathbf{1}))$ and $(\mu_{1}(\mathbf{0}),\ldots,\mu_{n}(\mathbf{0}))$ over the constraint defined by Assumption 2 (i); see Remark 10 in Appendix C.13 for details. Nevertheless, we can show that the equivalence holds up to a constant.
Theorem 6.4.
Under the conditions in Theorem 6.3, with $w_{i}\leq\bar{w}$ and $\|w\|_{1}\geq\underline{w}n$ for some constants $0<\underline{w}<\bar{w}<\infty$ that do not vary with $n$,
$1/4\leq\liminf_{K_{n}\rightarrow\infty}\frac{\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}[(\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\tau_{n,\mu})^{2}]}{\mathcal{B}_{n}^{*}(\mathcal{C}_{n},w,1)}\leq\limsup_{K_{n}\rightarrow\infty}\frac{\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}[(\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\tau_{n,\mu})^{2}]}{\mathcal{B}_{n}^{*}(\mathcal{C}_{n},w,1)}\leq 1$
Proof of Theorem 6.3.
See Appendix C.13. ∎
Theorem 6.4 justifies that the objective function $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},w,1)$ is a reasonable proxy for the worst-case MSE – its optimal solution yields a clustering whose worst-case MSE is no more than four times the optimal worst-case MSE.
As with Section 4, to obtain the frontier that trades off the worst-case bias and variance, we can remove the square in the second term and optimize the following proxy objective for a range of $\xi_{n}$:
$\frac{\xi_{n}}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}+\sum_{i=1}^{n}\frac{w_{i}}{||w||_{1}}\frac{1}{|\mathcal{N}_{i}|}\Big{|}\mathcal{N}_{i}\bigcap\Big{\{}j:c(i)\neq c(j)\Big{\}}\Big{|},$ (20)
@@
$\max_{\mathbf{X}(K)}\text{tr}(\mathbf{L}_{\xi_{n}}^{w}\mathbf{X}(K)),\quad\text{s.t.}\quad\text{diag}(\mathbf{X}(K))=\mathbf{1}_{n},\,\,\mathbf{X}(K)\succeq 0,\quad\mathbf{L}_{\xi_{n}}^{w}=||w||_{1}\mathbf{L}^{w}-\xi_{n}ww^{\top},$ (21)
@@
Algorithm 2 presents details. In summary, Theorem 6.3 shows how our insights in the characterization of the worst-case bias and variance more broadly apply to other estimators.
@@
1:Adjacency matrix $\mathbf{A}$, K, $\bar{K}$ smallest and largest number of clusters, $\xi_{n}$, number of iterations $T$
@@ 6.4 Saturation designs
denoting a difference between treated and control units, reweighted by the individual treatment probability, similar to the difference in means estimator studied in previous sections.
Theorem 6.5.
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}\Big{[}\Big{(}\mathbb{E}[\hat{\tau}_{n}^{q}(\mathcal{C}_{n})]-\hat{\tau}_{n}\Big{)}^{2}\Big{]}=$ $(q_{1}-q_{0})^{2}\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}[\bar{\psi}+o(1)],$ (23)
@@
Proof of Theorem 6.5.
See Appendix C.14. ∎
Theorem 6.5 characterizes the worst-case bias, variance, and MSE for estimating the global treatment effect under a saturation design. The bias depends on an additional bias component with respect to Lemma 3.1 that arises from possibly assigning individuals in the same cluster to different treatments. The bias under a saturation design increases because individuals in the same clusters are not all assigned to the same treatment status. On the other hand, the variance decreases because treatments are not necessarily perfectly correlated in the same cluster. For given treatment probabilities $(q_{1},q_{0})$ with $q_{1}+q_{0}=1$, and given clustering algorithms, researchers can compare such algorithms by studying their worst-case bias, variance, or MSE. In practice, if researchers are mostly concerned about the bias of the estimated overall treatment effects, cluster experiments are recommended.
In conclusion, Theorem 6.5 shows how our characterization of the bias and variance can be extended to more complex designs.
@@
denoting the contrast between potential outcomes under treatment when switching the treatment status of their friends from one to zero. The estimator compares average outcomes between treated units in clusters with larger and lower treatment probability. It is easy to show that the tools developed in this paper allow us to characterize the bias and variance of the estimator in Equation (24) relative to the estimand in the same equation. In particular, results similar to those in Theorem 6.5 for the overall treatment effect can also apply to other estimands (spillover effects in Equation 24, direct effects, and other forms of spillover effects) in the presence of saturation experiments discussed in the current subsection. ∎
7 Recommendations for practice
We conclude with a summary of our method and explicit recommendations for practice.
Choice of the adjacency matrix
The choice of the adjacency matrix must depend on researchers’ prior knowledge of which dimension spillovers propagate over, as, for example, in Cai et al. (2015). Our framework directly extends to settings where researchers only have some prior distribution on the network as in Breza et al. (2020). In this case, researchers can compute the expected bias and variance over the network distribution by taking the expected value of the expressions of the bias presented in Section 3 (see Remark 7).
Choosing the range of magnitude for the spillover effects
Similarly to standard power analysis (e.g. Baird et al., 2018), our method with $\lambda=1$ also depends on the choice of $\xi_{n}=(\bar{\phi}_{n}^{2}/\bar{\psi})^{-1}$ (the size of spillover effects $\bar{\phi}_{n}$ relative to the outcomes’ largest squared deviation $\bar{\psi}$). In our empirical applications, we observe that certain clustering algorithms uniformly outperform others for a large range of values of $\xi_{n}$, suggesting that studying clustering over ranges of $\xi_{n}$ can be informative.
To choose such a range of values, our recommended choice for $\bar{\psi}$ is to consider values $\bar{\psi}=c\bar{\sigma}^{2},c\in[1,4]$ and $\bar{\sigma}^{2}$ is the baseline variance of the residuals after removing the covariate adjustment, and $c$ is a constant between one to four. [^16] To gain further intuition, let $\lambda=1$. Let $\bar{\mu}_{i}$ be a prediction for $\mu_{i}(\mathbf{0})$ as in Remark 4 and consider an estimator as in Equation (4). When using regression adjusted estimators, we have $\bar{\psi}=\Big{(}\mu_{i}(\mathbf{1})-\bar{\mu}_{i}+\mu_{i}(\mathbf{0})-\bar{\mu}_{i}\Big{)}^{2}$. By approximating $(\mu_{i}(\mathbf{d})-\bar{\mu}_{i})^{2}\approx\bar{\sigma}^{2}$, we can write $\bar{\psi}\leq 4\bar{\sigma}^{2}$. Algorithm 3 presents details.
The choice of $\bar{\phi}_{n}$ can be based on spillover effects observed in previous experiments, in the spirit of minimum detectable effects (Baird et al., 2018), or using a priori bound on the magnitude of the spillover effects.
Choosing between a Cluster or Bernoulli design
Given the range of values of $\xi_{n}$, and the choice of the network, Theorem 3.6 provides a rule of thumb to choose between a cluster or Bernoulli design. Table 1 suggests a rule of thumb $\bar{\phi}_{n}\sqrt{K_{n}}>2.3$ for outcomes bounded between minus one and one. Figure 2 provides a wider and specific range of values of the minimum spillover effects to run cluster experiments.
Choosing the optimal clusters
Finally, once researchers decide to run a clustering experiment, Algorithm 1 provides an explicit algorithm to estimate the optimal clusters via semi-definite programming, choosing both the clustering and the number of clusters.
Algorithm 3 Practical choice of range of tuning parameter $\xi_{n}$
1:Baseline covariates $Z_{1},\cdots,Z_{n}$
2:Estimate $\bar{\mu}_{i}$ by regressing baseline outcomes on covariates $Z$
3:Estimate $\bar{\sigma}^{2}$ the variance of the residuals from this regression
4:Consider the range of values $\bar{\psi}\in[\bar{\sigma}^{2},4\bar{\sigma}^{2}]$
5:if past experiments are available then
6:     Use $\bar{\phi}_{n}$ from previous experiments
7:else
8:     Consider a range of values of plausible spillover effects
9:end if
Appendix A Notation
Table 4: Notation.
Description Mathematical Formulation
Spillovers’ effects size $\bar{\phi}_{n}$
Neighbors $\mathcal{N}_{i}$
Maximum degree $\mathcal{N}_{n,\max}=\max_{i}|\mathcal{N}_{i}|$
Treatments $D_{1},\cdots,D_{n},D_{i}\in\{0,1\}$
Cluster’s partition and size $\mathcal{C}_{n}$
Number of clusters $K_{n}$
Cluster’s groups $c(i)\in\{1,\cdots,K_{n}\}$
Bernoulli design $\mathcal{C}_{\mathrm{B},n}$
Appendix B Endogenous peer effects
Following models in Manski (1993), let
$Y_{i}(\mathbf{d})=\alpha+\beta_{n}\mathbf{d}_{i}+\kappa_{n}\frac{\sum_{j\neq i}\mathbf{A}_{i,j}\mathbf{d}_{j}}{\sum_{j\neq i}\mathbf{A}_{i,j}}+\gamma_{n}\frac{\sum_{j\neq i}\mathbf{A}_{i,j}Y_{i}(\mathbf{d})}{\sum_{j\neq i}\mathbf{A}_{i,j}}+\varepsilon_{i},\quad\beta_{n},\kappa_{n},\gamma_{n}\in[-\bar{\phi}_{n},\bar{\phi}_{n}],\bar{\phi}_{n}\in(-1,1),$

### v3

@@ B.2 Relaxing the ex-ante symmetry
The potential outcomes $\mu=(\mu_{1},\cdots,\mu_{n})\in\mathcal{M}=\otimes_{i=1}^{n}\mathcal{M}_{i}$ where $\mathcal{M}_{i}$ is the set such that $\mu_{i}(1,\cdot)\in\mathcal{M}_{1,i},\mu_{i}(0,\cdot)\in\mathcal{M}_{0,i}$. For all $i\in\{1,\cdots,n\}$, $d\in\{0,1\}$, the function classes $\mathcal{M}_{d,i}$ is the class of functions that include all functions $\mu_{i}(d,\cdot)$ satisfying the following two conditions
@@
    $\mu_{i}(d,\cdot)\in[\psi_{Li},\psi_{Ri}]$ for some $-\infty<\psi_{Li}<\psi_{Ri}<\infty$;
@@
    $\Big{|}\mu_{i}(0,\mathbf{d})-\mu_{i}(0,\mathbf{0})\Big{|}$ $\leq\bar{\phi}_{n}\frac{\alpha_{i}}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}\mathbf{d}_{k},$ (16)
    $\Big{|}\mu_{i}(1,\mathbf{d})-\mu_{i}(1,\mathbf{1})\Big{|}$ $\leq\bar{\phi}_{n}\frac{\alpha_{i}}{|\mathcal{N}_{i}|}\sum_{k\in\mathcal{N}_{i}}\Big{(}1-\mathbf{d}_{k}\Big{)},$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}(\mathcal{C}_{n})-\mathbb{E}[\hat{\tau}_{n}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\left\{\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}\right\}(1+\mathcal{O}(\varepsilon)),$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}[(\hat{\tau}_{n}(\mathcal{C}_{n})-\tau_{n,\mu})^{2}]=\left\{\bar{\phi}_{n}^{2}b_{n,\alpha}(\mathcal{C}_{n})^{2}+\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}\left(\frac{1}{n_{k}}\sum_{i\in c_{k}}\bar{\psi}_{i}\right)^{2}\right\}(1+\mathcal{O}(\varepsilon)).$
Proof of Theorem B.2.
See Appendix E.4. ∎
It is easy to show that similarly to what discussed in the main text, the optimization program admits a semi-definite relaxation that depends on the values of $\alpha$ and $\bar{\psi}$.
B.3 Trimming estimators
@@
$\sup_{\mu\in\mathcal{M}}\Big{|}\tau_{n,\mu}-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{|}=\left(\bar{\phi}_{n}b_{n}^{w}(\mathcal{C}_{n})+(\psi_{R}-\psi_{L})\sum_{i-1}^{n}\Big{|}\frac{w_{i}}{\|w\|_{1}}-\frac{1}{n}\Big{|}\right)\cdot(1+\mathcal{O}(\varepsilon))$
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}\Big{[}\Big{(}\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\mathbb{E}_{\mu}[\hat{\tau}_{n}^{w}(\mathcal{C}_{n})]\Big{)}^{2}\Big{]}=\frac{1}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}\cdot\bar{\psi}(1+\mathcal{O}(\varepsilon)).$
Proof of Theorem B.3.
See Appendix E.5. ∎
Theorem B.3 characterizes the worst-case bias and variance of the trimming estimator. Following the discussion for the difference in means estimator, this result suggests a weighted objective function
@@
For the difference in means estimator, the objective function $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},\lambda)$ defined in Theorem 3.3 can be interpreted as the worst-case MSE, up to $1+\mathcal{O}(\varepsilon)$ terms, as implied by Theorem 3.3. Unfortunately, this exact equivalence no longer holds for general trimming estimators because the worst-case bias and variance are achieved by different configurations of potential outcomes. In fact, we can show that the worst-case MSE is given by the maximum of a quadractic function in the vectors $(\mu_{1}(\mathbf{1}),\ldots,\mu_{n}(\mathbf{1}))$ and $(\mu_{1}(\mathbf{0}),\ldots,\mu_{n}(\mathbf{0}))$ over the constraint defined by Assumption 2 (i); see Remark 11 in Appendix E.6 for details. Nevertheless, we can show that the equivalence holds up to a constant.
Theorem B.4.
Under the conditions in Theorem B.3, with $w_{i}\leq\bar{w}$ and $\|w\|_{1}\geq\underline{w}n$ for some constants $0<\underline{w}<\bar{w}<\infty$ that do not vary with $n$,
$1/4\leq\liminf_{K_{n}\rightarrow\infty}\frac{\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}[(\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\tau_{n,\mu})^{2}]}{\mathcal{B}_{n}^{*}(\mathcal{C}_{n},w,1)(1+\mathcal{O}(\varepsilon))}\leq\limsup_{K_{n}\rightarrow\infty}\frac{\sup_{\mu\in\mathcal{M}}\mathbb{E}_{\mu}[(\hat{\tau}_{n}^{w}(\mathcal{C}_{n})-\tau_{n,\mu})^{2}]}{\mathcal{B}_{n}^{*}(\mathcal{C}_{n},w,1)(1+\mathcal{O}(\varepsilon))}\leq 1$
Proof of Theorem B.3.
See Appendix E.6. ∎
Theorem B.4 justifies that the objective function $\mathcal{B}_{n}^{*}(\mathcal{C}_{n},w,1)$ is a reasonable proxy for the worst-case MSE – its optimal solution yields a clustering whose worst-case MSE is no more than four times the optimal worst-case MSE.
As with Section 4, to obtain the frontier that trades off the worst-case bias and variance, we can remove the square in the second term and optimize the following proxy objective for a range of $\xi$:
$\frac{\xi}{\|w\|_{1}^{2}}\sum_{k=1}^{K_{n}}\left(\sum_{i:c(i)=k}w_{i}\right)^{2}+\sum_{i=1}^{n}\frac{w_{i}}{||w||_{1}}\frac{1}{|\mathcal{N}_{i}|}\Big{|}\mathcal{N}_{i}\bigcap\Big{\{}j:c(i)\neq c(j)\Big{\}}\Big{|},$ (19)
@@
$\max_{\mathbf{X}(K)}\text{tr}(\mathbf{L}_{\xi}^{w}\mathbf{X}(K)),\quad\text{s.t.}\quad\text{diag}(\mathbf{X}(K))=\mathbf{1}_{n},\,\,\mathbf{X}(K)\succeq 0,\quad\mathbf{L}_{\xi}^{w}=||w||_{1}\mathbf{L}^{w}-\xi ww^{\top},$ (20)
@@
Algorithm 2 presents details. In summary, Theorem B.3 shows how our insights in the characterization of the worst-case bias and variance more broadly apply to other estimators.
@@
1:Adjacency matrix $\mathbf{A}$, K, $\bar{K}$ smallest and largest number of clusters, $\xi$, number of iterations $T$
@@ B.4 Saturation designs
denoting a difference between treated and control units, reweighted by the individual treatment probability, similar to the difference in means estimator studied in previous sections. Noe that here we focus on $\hat{\tau}_{n}(\mathcal{C}_{n})$ to estimate the global effect for simplicity, but in principle one could consider alternative estimands and estimators, see Remark 10.
Theorem B.5.
@@
$\sup_{\mu\in\mathcal{M}}\mathbb{E}\Big{[}\Big{(}\mathbb{E}[\hat{\tau}_{n}^{q}(\mathcal{C}_{n})]-\hat{\tau}_{n}\Big{)}^{2}\Big{]}=$ $(q_{1}-q_{0})^{2}\sum_{k=1}^{K_{n}}\frac{n_{k}^{2}}{n^{2}}[\bar{\psi}+\mathcal{O}(\varepsilon)],$ (22)
@@
Proof of Theorem B.5.
See Appendix E.7. ∎
Theorem B.5 characterizes the worst-case bias, variance, and MSE for estimating the global treatment effect under a saturation design. The bias depends on an additional bias component with respect to Lemma 3.1 that arises from possibly assigning individuals in the same cluster to different treatments. The bias under a saturation design increases because individuals in the same clusters are not all assigned to the same treatment status. On the other hand, the variance decreases because treatments are not necessarily perfectly correlated in the same cluster. For given treatment probabilities $(q_{1},q_{0})$ with $q_{1}+q_{0}=1$, and given clustering algorithms, researchers can compare such algorithms by studying their worst-case bias, variance, or MSE. In practice, if researchers are mostly concerned about the bias of the estimated overall treatment effects, cluster experiments are recommended.
In conclusion, Theorem B.5 shows how our characterization of the bias and variance can be extended to more complex designs.
@@
denoting the contrast between potential outcomes under treatment when switching the treatment status of their friends from one to zero. The estimator compares average outcomes between treated units in clusters with larger and lower treatment probability. It is easy to show that the tools developed in this paper allow us to characterize the bias and variance of the estimator in Equation (23) relative to the estimand in the same equation. In particular, results similar to those in Theorem B.5 for the overall treatment effect can also apply to other estimands (spillover effects in Equation 23, direct effects, and other forms of spillover effects) in the presence of saturation experiments discussed in the current subsection. ∎
B.5 Endogeneous peer effects
Our results can extend to higher order spillover effects as long as we can write $Y_{i}(\mathbf{d})=\mu_{i}(\mathbf{d}_{i},\mathbf{d}_{\mathcal{N}_{i}})+\mathcal{O}(h_{n}),$ for some $h_{n}\rightarrow 0$. Our results hold if $h_{n}=o(1/n)$, capturing the idea that first-order effects $\bar{\phi}_{n}$ are larger than second-order effects.
We provide an example where this holds. Following models in Manski (1993), let
$Y_{i}(\mathbf{d})=\alpha+\beta_{n}\mathbf{d}_{i}+\kappa_{n}\frac{\sum_{j\neq i}\mathbf{A}_{i,j}\mathbf{d}_{j}}{\sum_{j\neq i}\mathbf{A}_{i,j}}+\gamma_{n}\frac{\sum_{j\neq i}\mathbf{A}_{i,j}Y_{i}(\mathbf{d})}{\sum_{j\neq i}\mathbf{A}_{i,j}}+\nu_{i},\quad\beta_{n},\kappa_{n},\gamma_{n}\in[-\bar{\phi}_{n},\bar{\phi}_{n}],\bar{\phi}_{n}\in(-1,1),$